UARS Weekly Status Report

09 November 2001




*********************************************************************

                UARS WEEKLY STATUS REPORT - 09 NOV 2001
							
********************************************************************

***********
SOLSTICE
***********

SOLSTICE continues normal operations, taking solar data on all available
orbital tracking periods and stellar calibration observations whenever
possible. We continue to gather enough data to meet the SOLSTICE science
objectives during this period of reduced tape recorder operations.

***********
SUSIM
***********

SUSIM completed daily mid resolution and MgII solar scans
during periods of predicted UARS telemetry.

We are working on software components which will allow us to 
process the SUSIM data from level 1 to level 3 on the SUSIM RAC.

***********
ACRIM
***********

Nothing received.

***********
PEM
***********

***** ALL PEM COMMUNICATION SHOULD NOW BE THROUGH THE PEM2K:: NODE *****

The PEM boom instruments were activated on 2001303 at about 22 UT in
full operational mode.

Thanks to all who helped Sandee Jeffers with PACOR-A files.

* Beginning of power orbit day (PEM activation): array current >5 amps, array 
voltage >53.5 volts, and battery current >2.0 amps.  
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.

***********
HALOE 
***********

Nothing received.

***********
MLS
***********

Nothing received.

***********
HRDI
***********

Nothing received.

***********
WINDII
***********

Nothing received.

***********
MPG
***********

The MPG continues to serve as the point-of-contact
for UARS instrument operations to support re-engineering
of the UARS Mission Operations Concept to develop 
capabilities for the Science Traceability Mission.

Daily processing of instrument Daily Activity Plans 
continues with no problems.  

***********
CDHF
***********

Nothing received.
                  
***********
FOT
***********

                          REPORTING PERIOD

This report covers 27 October 2001 (Orbit 55367, GMT Day 300) through 02 
November 2001 (Orbit 55471, GMT Day 306).

                       SPACECRAFT OPERATION

The observatory is now in Reverse Flight (Northern Hemisphere viewing) and is 
performing nominally under two-battery operations.  The instrument operational 
changes were:


                  INSTRUMENT OPERATIONAL CHANGES


     Instrument     Time       Orbit     Comment
         SOLSTICE     300/2112     55379     SOLSTICE REBOOT
         HALOE        302/1458     55406     Power OFF
         PEM          303/2254     55426     Day Only to FULLTIME ops
         WINDII       305/1500     55451     WINDII ERROR FIX
         HRDI         306/1315     55465     Safe Mode


The beta angle ranged between the angles of -41.0 degrees and 65.0 degrees for 
this report period.  The beta angle is now increasing towards a maximum angle 
of -2 degrees on 26 November 2001 (DOY 330).

Spacecraft battery 1, 2, and 3 performance monitoring continues.  Battery 1 
remains OFF the charge relay.  Battery 2 and Battery 3 are currently 
maintaining greater than 23.8 V end-of-night (EON) load bus voltage.  Battery 
temperatures are stable with a temperature delta between Battery 2 and Battery 
3 of 0.6 and 0.9 degrees C.  Additional battery performance data for each 
battery on 27 October 2001 and 02 November 2001 are provided in Appendix D.

The clock error ranged from 10.2 to 16.7 msec during this report period.  
Clock rate adjustments are listed in Appendix C.

All instrument and subsystem engineers were advised on the initial excursion of 
ALERTS or Out-Of-Limits conditions detected by the FOT for this report.  A 
summary is listed below and the details of ALERT and Out-Of-Limits occurrences 
are listed in Appendix B.

     OUT-OF-LIMITS  =  03
     ALERTS         =  01

The following real-time command sheets were executed for the listed operational 
element(s) during this report period (see Appendix A).

     FOT            =  14
     SOLSTICE       =  1
     HALOE          =  1
     PEM            =  1
     WINDII         =  1
     HRDI           =  1


UARS Anomaly Reports initiated during this report period are listed in Appendix 
E and summarized below.

     FOT            =  02
     SOLSTICE       =  01
     WINDII         =  01
     ACRIM          =  01

One (1) Anomaly Report remained OPEN at the end of this report period (see 
Appendix E).

TRANSPONDER FREQUENCIES:
 Transponder A:  00/316  - 2287.496493
 Transponder B:  01/286  - 2287.498613
 
                      GROUND SYSTEM OPERATION


Five (5) operational support problems occurred this period.  Anomaly Reports 
are included in Appendix E.

The data loss calculations are no longer being supplied by Data Capture 
Facility (DCF). This service should resume soon.


                      OTHER SIGNIFICANT EVENTS

                      UARS SATELLITE OPERATIONS

The FOT continued to work on getting the TAC replacement system ready for 
operations.  Two urgent DRs were written against the system, which must be 
resolved before the TAC will be ready for operations. The FOT and software 
developers performed data analysis of dumps from the old and new TAC to 
determine the cause of the errors found in the new TAC.  The source of the 
errors is has yet to be found, but should be found soon.  It is important to 
resolve these issues quickly due to the tight schedule of the ground system 
replacement.

The FOT made preparations for the upcoming SRR for the Eclipse ground system.  
The goal of the SRR will be to make sure all of the requirement for the new 
system have been found and are well understood.


                            MEETINGS

The FOT attended a UARS Ground System Rehost Meeting on Thursday, November 01, 
2001
                     FUTURE SIGNIFICANT EVENTS

The next UARS Yaw slew is scheduled for 26 November 2001 (DOY 330).  This 
maneuver will be a Reverse to Forward Flight maneuver (Southern Hemisphere 
viewing).

Attachments:  Appendix A,  Real-time Commands Sheets
              Appendix B,  Out-Of-Limits & ALERT Occurrences
              Appendix C,  Clock & Solar Array Rate Corrections
              Appendix D,  UARS Battery Performance Data
              Appendix E,  UARS Anomaly Reports

                            APPENDIX A
                     REAL-TIME COMMAND SHEETS
                    Week ending 02 November 2001

ORBIT     AOS          NO.       SUBSYS     REASON
-----     ---          ---        ------    ------
55367     300/0035     01-648     NBTR      NBTR PBK
55371     300/0811     01-648               NOOPS D-300
55378     300/1830     01-649     NBTR      NOOPS D-301
55379     300/2112     01-650     SOLSTICE  SOLSTICE REBOOT
55382     301/0036     01-649               NBTR PLAYBACK
55386     301/0811     01-649     NBTR      NOOPS D-301
55399     302/0410     01-651     NBTR      NBTR PBK
55400     302/0638     01-651     NBTR      NOOPS D-302
55401     302/0812     01-652     OBC       Clk Adj 0.0 to 20.6 msec/day
55406     302/1458     01-654     HALOE     Power OFF
55408     302/1832     01-653     NBTR      NOOPS D-303
55413     303/0220     01-653     NBTR      NBTR PBK
55415     303/0630     01-653     NBTR      NOOPS D-303
55426     303/2254     01-655     PEM       Day Only to FULLTIME ops
55436     304/1459     01-656     OBC       Clk Adj 20.6 to 0.0 msec/day
55451     305/1500     01-657     WINDII    WINDII ERROR FIX
55458     306/0229     01-658     OBC       Clk Adj 0.0 to 20.6 msec/day
55460     306/0630     01-659     OBC       Clk Adj 20.6 to 41.2 msec/day
55465     306/1315     01-660     HRDI      Safe Mode
_______________________________________________________________________________

                             APPENDIX B
                   OUT-OF-LIMITS & ALERTS OCCURRENCES
                     Week ending 02 November 2001

The following Out-Of-Limits occurrences were detected:

MNEMONIC         STATE     DESCRIPTION              ORBITS
--------         -----     -----------              ------
HRPTSBARTEMP     Y-HI      HRDI PTS Barrel Temp     55455,56
HRPTSBARTEMP     Y-LO      HRDI PTS Barrel Temp     55457-63
HRPTSBARTEMP     R-LO      HRDI PTS Barrel Temp     55464


The following ALERT occurrences were detected:

MNEMONIC       DESCRIPTION                         ORBITS
--------       -----------                         ------
WIERRLIMIT     WINDII Software Error Code          55440-451

_______________________________________________________________________________

                             APPENDIX C
              CLOCK & SOLAR ARRAY RATE CORRECTIONS
                   Week ending 02 November 2001

                   CLOCK DRIFT RATE ADJUSTMENTS

ORBIT     TIME         CLOCK RATE CHANGE
-----     -----        -----------------
55401     302/0815     Clk Adj 0.0 to 20.6 msec/day
55436     304/1509     Clk Adj 20.6 to 0.0 msec/day
55458     306/0241     Clk Adj 0.0 to 20.6 msec/day
55460     306/0644     Clk Adj 20.6 to 41.2 msec/day
___________________________________________________________________________

                             APPENDIX D
                    UARS BATTERY PERFORMANCE DATA
                     Week ending 02 November 2001


     27 October 2001 - GMT Day 300 Beta = -41.0 deg, SA Pos = 269 deg

     DIFF V (mV)     TEMP (C)      CURR (amp)     EON V     AVE C/D   AVE
BAT  MAX/MIN         MAX/MIN       MAX/MIN        MIN       RATIO     DOD(%)
---  ----------      --------      ---------      -----     -------   -----
1    +00.0/-22.4     -3.05/-4.62   +0.0/+0.0      13.0      0.000     0.0
2    +11.2/-22.4     +4.51/+3.23   +27.2/-13.6    26.4      1.030     19.5
3    +22.4/-22.4     +3.87/+2.60   +26.4/-13.6    26.2      1.030     19.5


     02 November 2001 - GMT Day 306 Beta = -65.0 deg, SA Pos = 269 deg

     DIFF V (mV)     TEMP (C)      CURR (amp)     EON V     AVE C/D   AVE
BAT  MAX/MIN         MAX/MIN       MAX/MIN        MIN       RATIO     DOD(%)
---  ------------    ----------    -----------    -------   -------   ------
1    +16.8/+00.0     -3.68/-4.62   +0.0/+0.0      13.0      0.000     0.0
2    +16.8/-44.8     +5.15/+3.87   +26.8/-15.2    25.1      1.030     20.5
3    +22.4/-67.2     +4.51/+2.92   +26.4/-15.2    25.1      1.020     20.2


     V/T LEVEL CHANGES
ORB#           TIME          LVL FR-TO
----           ----          ---------
-no changes    V/T 5

_______________________________________________________________________________

                                  APPENDIX E
                      UARS ANOMALY REPORTS
                    Week ending 02 November 2001

                   NEW ANOMALY REPORTS GENERATED
AR NO.     SUBSYS     ORBIT     TIME        TTR     PROBLEM/DATA LOSS
------     ------     -----     ----        ---     -----------------
01-066     FOT        55378     300/1931Z   N/A     Exec Stat Buffer Error
01-067     SOLSTICE   55379     300/2112Z   N/A     Detector Tubes Off
01-068     WINDII     55440     304/2126Z   N/A     Error Code E11
01-069     FOT        55451     305/1500Z   14170   Generic Late Acquisition
01-070     ACRIM      55453     305/1931Z   N/A     ACRIM In "OFF" State


               ANOMALY CLOSURE INFORMATION RECEIVED
AR NO.     SUBSYS     ORBIT     TIME        TTR     PROBLEM/DATA LOSS
------     ------     -----     ----        ---     -----------------
01-063     FOT        see list  Oct 01      14170   Generic Late Acquisition
01-066     FOT        55378     300/1931Z   N/A     Exec Stat Buffer Error
01-067     SOLSTICE   55379     300/2112Z   N/A     Detector Tubes Off
01-068     WINDII     55440     304/2126Z   N/A     Error Code E11
01-070     ACRIM      55453     305/1931Z   N/A     ACRIM In "OFF" State

             ANOMALY INVESTIGATION REPORTS GENERATED
AR NO.     SUBSYS     ORBIT     TIME       TTR      PROBLEM/DATA LOSS
------     ------     -----     ----       ---      -----------------
-None this report period

                       OPEN ANOMALY REPORTS
AR NO.     SUBSYS     ORBIT     TIME       TTR      PROBLEM/DATA LOSS
------     ------     -----     ----       ---      -----------------
01-69     FOT         see list  Nov 01     14170    Generic Late Acquisition



	 	SEVNTFW01/44